A material has a heat capacity at constant volume C that is unchanging over a wide temperature
range. Two identical bodies made of this material, at initial temperature T and 9T, respectively, are
used to drive a heat engine. The maximum amount of work that can be obtained from this engine is
1
Expert's answer
2020-05-03T17:16:16-0400

"W=\eta Q=C(9T-T)\left(1-\frac{T}{9T}\right)=\frac{64}{9}CT"